1/3(2-5p)= -13 please help me solve this step by step thanks
Answer by midwood_trail(310)

1/3(2-5p)= -13
The first thing you want to do is remove the parentheses whenever you see one in any math question regardless of grade.
We can write 1/3(2-5p) as (2 - 5p)/3 and then equate it to -13.
It looks like this:
(2 - 5p)/3 = - 13
We now want to remove that ugly fraction on the left side.
To do so, we multiply both sides of the equation by the LCD, which is 3 in this case.
So, (2 - 5p)/3 times 3 = 2 - 5p.
-13 times 3 = -39.
We now have an equation without a fraction and it looks like this:
2 - 5p = -39
We now solve for P.
Subtracting 2 from both sides of the equation and we get:
-5p = -41
To find p, divide both sides by -5.
p = -41/-5
p = 41/5
If your teacher is looking for a decimal answer, then divide 41 by 5 and your answer for p will be 8.2
I hope that helps.